If the sodium concentration in blood plasma is 0.140 M, and Ksp for sodium urate is 5.76×10−8, what minimum concentration of urate would result in the precipitation of odium urate?
Answer #1

The equilibrium

C5H3N4O3Na ---------> C5H3N4O3-   +Na+

Ksp    = [C5H3N4O3-] [Na+]
let x moles/L sodium urate dissolve will produce x mole/L urate and x moles of Na+
but already some Na+ present at equilibrium
[Na+] = 0.140+x and [Urate] =x

Ksp    = [C5H3N4O3-] [Na+]

5.76*10-8    = x*(0140+x)

5.76*10-8 = 0.14x + x2

x2 + 0.14x-5.76*10-8 =0

     x= 4.1*10-7 M

[urate] = x = 4.1*10-7 M